# Vibe AIGC: A New Paradigm for Content Generation via Agentic Orchestration

### v1

@@ 3.1 Prevailing Architectures in Video Generation
The field of generative AI has progressed from static image synthesis to dynamic video generation, with foundational models now demonstrating significant advancements in text alignment, visual fidelity, motion plausibility, and realism Liu et al. (2025b). Current research mainly focuses on text-to-image (T2I) Rombach et al. (2022), text-to-video (T2V) Yin et al. (2025), and image-to-video (I2V) generation Xing et al. (2024), as these modalities allow for the collection of scalable, high-quality datasets Wang et al. (2025) curated through rigorous filtering and captioning processes Chen et al. (2025c).
The latest video generation paradigm is the latent diffusion model with a spacetime Transformer, supplanting earlier GAN Goodfellow et al. (2020) and VQ-VAE-based Van Den Oord et al. (2017) methods. This approach first compresses a video into a lower-dimensional latent space of “spacetime patches” and then employs a diffusion Transformer (DiT) Peebles and Xie (2023) to denoise these patches conditioned on a text prompt. This patch-based framework offers the flexibility to process videos of variable resolutions, durations, and aspect ratios within a single model.
For I2V generation, the dominant strategy is to adapt a pre-trained T2V model. For example, Stable Video Diffusion Blattmann et al. (2023) fine-tunes the Stable Diffusion model by incorporating temporal layers, enabling it to generate motion that is coherent with a given input frame. Wan compresses the first frame into VAE latents and concatenates them with the noise latent along the channel axis Wan et al. (2025), so that the starting frame would be exactly maintained. A critical challenge in this area is maintaining identity consistency, ensuring a subject’s appearance remains stable across frames; techniques like IP-Adapter Ye et al. (2023) and integrated memory mechanisms Yu et al. (2025) have shown promise in addressing this challenge.
However, a fundamental constraint is the immense computational cost of video data, which results in training datasets that are smaller in scale and conceptual breadth than those for LLMs. This creates a disparity between the model’s limited world knowledge and high user expectations for both semantic understanding and visual fidelity. Consequently, prompt engineering (PE) Liu and Chilton (2022) becomes essential to bridge this gap by aligning user queries with the model’s learned data distribution. While sophisticated prompting can unlock impressive reasoning capabilities Wiedemer et al. (2025), the generation process often remains stochastic and requires considerable trial-and-error, increasing the barrier to effective use.
@@ 3.2 Editing and Reference-Based Video Generation
Advancing beyond unconditional generation from a single prompt Team et al. (2025a), a significant research frontier is the development of methods for granular control over video generation. Reference-based generation Ku et al. (2024) aims to transfer specific attributes from a source to a newly generated video. For instance, style transfer applies the visual aesthetic of a reference image or video onto a target video. Similarly, subject-driven generation Team et al. (2025b), inspired by personalization techniques like DreamBooth Ruiz et al. (2023), first learns a unique identifier for a subject from reference images. A diffusion model, conditioned on this identifier and a motion sequence, can then generate a new video of that subject. A primary obstacle for all reference-based tasks is the difficulty in acquiring training data; it requires meticulously constructed pairs of reference and target videos with guaranteed correspondence and quality. A common failure mode is “content leakage”, where unintended attributes from the reference are mistakenly rendered in the output. Such artifacts, often stemming from inherent model limitations, are difficult for users to mitigate through prompt engineering alone.
Video editing Jiang et al. (2025b); Wei et al. (2025a) focuses on modifying an existing video according to user instructions. The most direct approach is text-guided video editing He et al. (2025), which requires a model to comprehend textual commands. Representative methods, such as Senorita Zi et al. (2025), focus on object-level editing, including object addition, deletion, or modification. To improve precision, some works utilize masks Cai et al. (2025) to define the region for editing, tasking the model with in-painting the masked area conditioned on the text prompt. Similar to reference-based methods, training editing models is constrained by data availability, as paired “before” and “after” real-world videos hardly exist. This reliance on synthetic training data can lead to artifacts like pixel misalignment and unintended subject alterations.
@@ 3.3 Unified Architectures for Video Understanding and Generation
A forward-looking research direction is the development of single, unified models capable of both understanding and generation Cui et al. (2025); Lin et al. (2025). This approach treats video not as a specialized data type but as another modality, akin to text or audio, to be processed within a large-scale, multi-modal architecture Teng et al. (2025).
The mainstream unification tends to be achieved by adopting core architectural principles from LLMs. The critical first step is the discretization of continuous video data into a sequence of tokens. This is typically accomplished using a VQ-VAE or a similar network, which learns a codebook of visual patterns and compresses video patches into discrete codes. Once tokenized, the video sequence, now represented as a series of integers, can be processed by a standard Transformer architecture alongside tokens from other modalities like text. Within this framework, diverse tasks are framed as a unified next-token prediction problem. Consequently, tasks such as T2V, video captioning, video prediction, and video editing can be handled by a single model. However, this approach still faces challenges related to data quality and the need for exhaustive task enumeration. Empirically, these unified models still performs struggling on the fundamental T2V and I2V tasks, lagging behind specialized DiT models in terms of both generation fidelity and semantic alignment Huang et al. (2024); Ghosh et al. (2023).
@@ 3.4 Analysis of Realworld Workflows
Video generation models are increasingly adopted in various application domains, including animation, motion comics, news broadcasting, and film post-production, where they can simplify traditional content creation and reduce costs. However, due to the inherent limitations of current models—such as the stochastic nature of their outputs and unpredictable capabilities—users must devise intricate, multi-step workflows to achieve desired results. For instance, a creator producing a short-form drama might first develop a storyboard, generate a keyframe for each shot, and then iteratively refine prompts to produce short video segments, typically 5 or 10 seconds in length Wan et al. (2025); Kong et al. (2024), as constrained by the video generation service. These segments often suffer from artifacts Ye et al. (2025) like color discrepancies, inconsistent character identity, poor realism, and incorrect pacing or lip synchronization Hu et al. (2025) (even with audio-video joint generation Huang et al. (2025)). Consequently, significant manual post-processing is required to edit and assemble these clips into a coherent whole, followed by audio dubbing and video super-resolution Du et al. (2025). This process is inefficient, with substantial time and computational resources expended on failed generation attempts, making the outcome heavily reliant on trial and error. While users can produce a viable prototype, a more intelligent system is needed: one that can comprehend the user’s high-level creative intent, automatically decompose the task, and orchestrate the use of various tools, rather than depending solely on the combination of raw model outputs and manual prompt engineering.
@@ 4.1 Vibe AI-Generated Text Content
The inception of any creative work requires a deep understanding of the underlying subject matter. Traditional AIGC workflows often suffer from “contextual shallowness” due to the static nature of pre-trained knowledge. Recent advancements in autonomous reasoning, most notably OpenAI’s Deep Research OpenAI (2025), have demonstrated the potential for long-horizon information synthesis. By employing LLM-based agents to perform multi-step web searches, cross-verify disparate sources, and synthesize comprehensive knowledge bases, these systems build a robust semantic foundation before content generation begins. This “thinking-before-creating” paradigm is essential for Vibe AIGC, as it ensures that the generated content is grounded in a sophisticated aesthetic and factual context, moving beyond the limitations of single-model-based generation.
@@
In the traditional scholarly promotion workflow, researchers are often reduced to “manual dispatchers”. Even with the assistance of LLMs, the process remains model-centric and fragmented: authors must manually juggle multiple LLM interfaces for summarization, download and crop figures from PDFs, and painstakingly rewrite content to satisfy the distinct technical constraints and “vibes” of various social platforms. Thus, researchers introduced AutoPR Chen et al. (2025b) in Figure 2, a novel task that formalizes the transformation of research papers into accurate, engaging public content, and proposed a collaborative multi-agent system, comprising Logical Draft, Visual Analysis, and Textual Enriching agents.
@@ 4.2 Vibe AI-Generated Image Content
In professional graphic design, the primary challenge is the precise control over spatial layout and typography. Poster Copilot Wei et al. (2025b) in Figure 3 explored agentic layout reasoning and controllable editing. Unlike black-box generators, Poster Copilot functions as a design partner that translates abstract “Vibe” instructions into concrete design parameters such as geometric composition, color palettes, and layer hierarchies. By incorporating a feedback loop with human-in-the-loop editing, this system demonstrates how agents can bridge the gap between vague human aesthetic preferences and the rigid technical requirements of professional design.
@@ 4.3 Vibe AI-Generated Video Content
To tackle the complexities of music video (MV) creation, where visuals must align with rhythm, lyrics, and emotional arcs, researchers developed AutoMV Tang et al. (2025). In Figure 4. AutoMV represents a shift toward a collaborative multi-agent pipeline. The system employs a Screenwriter Agent to draft narrative scripts based on musical attributes (e.g., beats and structure) and a Director Agent to manage a shared Character Bank and coordinate with various video generation tools. This framework ensures that different segments of a full-length song remain visually and stylistically consistent. The success of AutoMV underscores the necessity of a modular, role-playing agentic structure in managing the high-level intent of a creative project.
In addition to the aforementioned works, there are numerous studies such as MotivGraph-SoIQ Lei et al. (2025), VideoAgent Wang et al. (2024), HollywoodTown Wei et al. (2025c), and LVAS-Agent Zhang et al. (2025). All preliminary efforts reveal a clear trajectory: the future of AIGC lies in orchestrating specialized agents capable of reasoning, planning, and maintaining long-term consistency. However, these systems remain largely fragmented within their respective domains. This observation serves as the primary catalyst for Vibe AIGC, which seeks to unify these agentic capabilities into a cohesive, intent-driven ecosystem.
@@ 5.1 Top-Level Architecture
At the apex of this architecture, the Meta Planner serves as the primary commander at the human-computer interface. Rather than executing generation tasks, it is responsible for receiving natural language and translating it into global system scheduling Jiang et al. (2025a). This process transcends simple keyword matching, employing reasoning-based dynamic construction Xiong et al. (2025). To ensure professional precision, the Meta Planner interacts deeply with an external domain-specific knowledge, storing professional skills, experiential knowledge, and a comprehensive registry of supported algorithmic workflows. For instance, when the Planner receives a request for an “oppressive atmosphere”, it queries the knowledge base to deconstruct this abstract “vibe" into specific engineering constraints—such as low-key lighting, close-ups, and low-saturation filters—thereby mitigating hallucinations and mediocrity common in general LLMs Li et al. (2025).
Regarding execution logic, the high-level design adopts a hierarchical orchestration strategy, mapping complex generation tasks through top-down layers of abstraction. The Meta Planner first generates a macro-level SOP blueprint at the creative layer, then propagates this logic to the algorithmic layer to automatically derive and configure the workflow graph structure Qiu et al. (2025). This hierarchical design ensures the system can both comprehend the “director’s vision” at a macro level and precisely control “technician operations” at a micro level. In essence, the top-level design of Vibe AIGC is not a static toolkit but a dynamic decision flow driven by the Meta Planner: it perceives the user’s “Vibe" in real-time, disambiguates intent via expert knowledge, and ultimately grows a precise, executable workflow from the top down Xiong et al. (2025).
@@ 5.2 Meta Planner
As the cognitive core of the Vibe AIGC architecture, the Meta Planner assumes the critical responsibility of translating natural language intent into an executable system architecture. Departing from the paradigm of traditional Large Language Models (LLMs) that function merely as text generators or simple routers, the Meta Planner is engineered as a “System Architect" endowed with high-level reasoning capabilities Goebel and Zips (2025). Positioned at the forefront of human-computer interaction, it directly interfaces with the user’s natural language input. Its primary function is not the immediate generation of content, but rather deep intent parsing and task decomposition. Upon receiving a user’s “Commander Instructions"—which are often ambiguous and unstructured—the Planner identifies explicit functional requirements while simultaneously capturing latent "Vibe" signals, such as style, mood, and rhythm. Through multi-hop reasoning, it converts these signals into internal logical representations, thereby triggering the entire generative engine.
@@ 5.3 Intent Understanding
The reasoning depth of the Meta Planner stems from its tight synergy with the domain-expert knowledge. This process is designed to flesh out sparse, subjective user instructions into actionable and objective creative schemes, thereby addressing intent sparsity Fagnoni et al. (2025).
The Planner begins by querying the creative expert knowledge modules, which encapsulate a vast array of multi-disciplinary expertise. For instance, when a user provides the instruction for “a Hitchcockian suspenseful opening," the Planner does not merely treat it as a text prompt. Instead, it leverages the knowledge base to deconstruct the abstract concept of “suspense" into a series of precise creative constraints: visually, it mandates "dolly zoom" camera movements and “high-contrast lighting"; auditorily, it requires "dissonant intervals" in the score; and narratively, it dictates an editing rhythm based on “information asymmetry." Through this process, the Planner externalizes implicit knowledge, transforming the user’s subjective aesthetic intuition into objective, concrete creative scripts. Consequently, this mitigates the issues of “averaging" (mediocrity) or "hallucination" often found in traditional AIGC tools due to comprehension gaps Shi et al. (2025).

### v2

@@ 3.1 Prevailing Architectures in Video Generation
The field of generative AI has progressed from static image synthesis to dynamic video generation, with foundational models now demonstrating significant advancements in text alignment, visual fidelity, motion plausibility, and realism (Liu et al., 2025b). Current research mainly focuses on text-to-image (T2I) (Rombach et al., 2022), text-to-video (T2V) (Yin et al., 2025), and image-to-video (I2V) generation (Xing et al., 2024), as these modalities allow for the collection of scalable, high-quality datasets (Wang et al., 2025) curated through rigorous filtering and captioning processes (Chen et al., 2025c).
The latest video generation paradigm is the latent diffusion model with a spacetime Transformer, supplanting earlier GAN (Goodfellow et al., 2020) and VQ-VAE-based (Van Den Oord et al., 2017) methods. This approach first compresses a video into a lower-dimensional latent space of “spacetime patches” and then employs a diffusion Transformer (DiT) (Peebles and Xie, 2023) to denoise these patches conditioned on a text prompt. This patch-based framework offers the flexibility to process videos of variable resolutions, durations, and aspect ratios within a single model.
For I2V generation, the dominant strategy is to adapt a pre-trained T2V model. For example, Stable Video Diffusion (Blattmann et al., 2023) fine-tunes the Stable Diffusion model by incorporating temporal layers, enabling it to generate motion that is coherent with a given input frame. Wan compresses the first frame into VAE latents and concatenates them with the noise latent along the channel axis (Wan et al., 2025), so that the starting frame would be exactly maintained. A critical challenge in this area is maintaining identity consistency, ensuring a subject’s appearance remains stable across frames; techniques like IP-Adapter (Ye et al., 2023) and integrated memory mechanisms (Yu et al., 2025) have shown promise in addressing this challenge.
However, a fundamental constraint is the immense computational cost of video data, which results in training datasets that are smaller in scale and conceptual breadth than those for LLMs. This creates a disparity between the model’s limited world knowledge and high user expectations for both semantic understanding and visual fidelity. Consequently, prompt engineering (PE) (Liu and Chilton, 2022) becomes essential to bridge this gap by aligning user queries with the model’s learned data distribution. While sophisticated prompting can unlock impressive reasoning capabilities (Wiedemer et al., 2025), the generation process often remains stochastic and requires considerable trial-and-error, increasing the barrier to effective use.
@@ 3.2 Editing and Reference-Based Video Generation
Advancing beyond unconditional generation from a single prompt (Team et al., 2025a), a significant research frontier is the development of methods for granular control over video generation. Reference-based generation (Ku et al., 2024) aims to transfer specific attributes from a source to a newly generated video. For instance, style transfer applies the visual aesthetic of a reference image or video onto a target video. Similarly, subject-driven generation (Team et al., 2025b), inspired by personalization techniques like DreamBooth (Ruiz et al., 2023), first learns a unique identifier for a subject from reference images. A diffusion model, conditioned on this identifier and a motion sequence, can then generate a new video of that subject. A primary obstacle for all reference-based tasks is the difficulty in acquiring training data; it requires meticulously constructed pairs of reference and target videos with guaranteed correspondence and quality. A common failure mode is “content leakage”, where unintended attributes from the reference are mistakenly rendered in the output. Such artifacts, often stemming from inherent model limitations, are difficult for users to mitigate through prompt engineering alone.
Video editing (Jiang et al., 2025b; Wei et al., 2025a) focuses on modifying an existing video according to user instructions. The most direct approach is text-guided video editing (He et al., 2025), which requires a model to comprehend textual commands. Representative methods, such as Senorita (Zi et al., 2025), focus on object-level editing, including object addition, deletion, or modification. To improve precision, some works utilize masks (Cai et al., 2025) to define the region for editing, tasking the model with in-painting the masked area conditioned on the text prompt. Similar to reference-based methods, training editing models is constrained by data availability, as paired “before” and “after” real-world videos hardly exist. This reliance on synthetic training data can lead to artifacts like pixel misalignment and unintended subject alterations.
@@ 3.3 Unified Architectures for Video Understanding and Generation
A forward-looking research direction is the development of single, unified models capable of both understanding and generation (Cui et al., 2025; Lin et al., 2025). This approach treats video not as a specialized data type but as another modality, akin to text or audio, to be processed within a large-scale, multi-modal architecture (Teng et al., 2025).
The mainstream unification tends to be achieved by adopting core architectural principles from LLMs. The critical first step is the discretization of continuous video data into a sequence of tokens. This is typically accomplished using a VQ-VAE or a similar network, which learns a codebook of visual patterns and compresses video patches into discrete codes. Once tokenized, the video sequence, now represented as a series of integers, can be processed by a standard Transformer architecture alongside tokens from other modalities like text. Within this framework, diverse tasks are framed as a unified next-token prediction problem. Consequently, tasks such as T2V, video captioning, video prediction, and video editing can be handled by a single model. However, this approach still faces challenges related to data quality and the need for exhaustive task enumeration. Empirically, these unified models still performs struggling on the fundamental T2V and I2V tasks, lagging behind specialized DiT models in terms of both generation fidelity and semantic alignment (Huang et al., 2024; Ghosh et al., 2023).
@@ 3.4 Analysis of Realworld Workflows
Video generation models are increasingly adopted in various application domains, including animation, motion comics, news broadcasting, and film post-production, where they can simplify traditional content creation and reduce costs. However, due to the inherent limitations of current models—such as the stochastic nature of their outputs and unpredictable capabilities—users must devise intricate, multi-step workflows to achieve desired results. For instance, a creator producing a short-form drama might first develop a storyboard, generate a keyframe for each shot, and then iteratively refine prompts to produce short video segments, typically 5 or 10 seconds in length (Wan et al., 2025; Kong et al., 2024), as constrained by the video generation service. These segments often suffer from artifacts (Ye et al., 2025) like color discrepancies, inconsistent character identity, poor realism, and incorrect pacing or lip synchronization (Hu et al., 2025) (even with audio-video joint generation (Huang et al., 2025)). Consequently, significant manual post-processing is required to edit and assemble these clips into a coherent whole, followed by audio dubbing and video super-resolution (Du et al., 2025). This process is inefficient, with substantial time and computational resources expended on failed generation attempts, making the outcome heavily reliant on trial and error. While users can produce a viable prototype, a more intelligent system is needed: one that can comprehend the user’s high-level creative intent, automatically decompose the task, and orchestrate the use of various tools, rather than depending solely on the combination of raw model outputs and manual prompt engineering.
@@ 4.1 Vibe AI-Generated Text Content
The inception of any creative work requires a deep understanding of the underlying subject matter. Traditional AIGC workflows often suffer from “contextual shallowness” due to the static nature of pre-trained knowledge. Recent advancements in autonomous reasoning, most notably OpenAI’s Deep Research  (OpenAI, 2025), have demonstrated the potential for long-horizon information synthesis. By employing LLM-based agents to perform multi-step web searches, cross-verify disparate sources, and synthesize comprehensive knowledge bases, these systems build a robust semantic foundation before content generation begins. This “thinking-before-creating” paradigm is essential for Vibe AIGC, as it ensures that the generated content is grounded in a sophisticated aesthetic and factual context, moving beyond the limitations of single-model-based generation.
@@
In the traditional scholarly promotion workflow, researchers are often reduced to “manual dispatchers”. Even with the assistance of LLMs, the process remains model-centric and fragmented: authors must manually juggle multiple LLM interfaces for summarization, download and crop figures from PDFs, and painstakingly rewrite content to satisfy the distinct technical constraints and “vibes” of various social platforms. Thus, researchers introduced AutoPR  (Chen et al., 2025b) in Figure 2, a novel task that formalizes the transformation of research papers into accurate, engaging public content, and proposed a collaborative multi-agent system, comprising Logical Draft, Visual Analysis, and Textual Enriching agents.
@@ 4.2 Vibe AI-Generated Image Content
In professional graphic design, the primary challenge is the precise control over spatial layout and typography. Poster Copilot  (Wei et al., 2025b) in Figure 3 explored agentic layout reasoning and controllable editing. Unlike black-box generators, Poster Copilot functions as a design partner that translates abstract “Vibe” instructions into concrete design parameters such as geometric composition, color palettes, and layer hierarchies. By incorporating a feedback loop with human-in-the-loop editing, this system demonstrates how agents can bridge the gap between vague human aesthetic preferences and the rigid technical requirements of professional design.
@@ 4.3 Vibe AI-Generated Video Content
To tackle the complexities of music video (MV) creation, where visuals must align with rhythm, lyrics, and emotional arcs, researchers developed AutoMV  (Tang et al., 2025). In Figure 4. AutoMV represents a shift toward a collaborative multi-agent pipeline. The system employs a Screenwriter Agent to draft narrative scripts based on musical attributes (e.g., beats and structure) and a Director Agent to manage a shared Character Bank and coordinate with various video generation tools. This framework ensures that different segments of a full-length song remain visually and stylistically consistent. The success of AutoMV underscores the necessity of a modular, role-playing agentic structure in managing the high-level intent of a creative project.
In addition to the aforementioned works, there are numerous studies such as MotivGraph-SoIQ (Lei et al., 2025), VideoAgent (Wang et al., 2024), HollywoodTown (Wei et al., 2025c), and LVAS-Agent (Zhang et al., 2025). All preliminary efforts reveal a clear trajectory: the future of AIGC lies in orchestrating specialized agents capable of reasoning, planning, and maintaining long-term consistency. However, these systems remain largely fragmented within their respective domains. This observation serves as the primary catalyst for Vibe AIGC, which seeks to unify these agentic capabilities into a cohesive, intent-driven ecosystem.
@@ 5.1 Top-Level Architecture
At the apex of this architecture, the Meta Planner serves as the primary commander at the human-computer interface. Rather than executing generation tasks, it is responsible for receiving natural language and translating it into global system scheduling (Jiang et al., 2025a). This process transcends simple keyword matching, employing reasoning-based dynamic construction (Xiong et al., 2025). To ensure professional precision, the Meta Planner interacts deeply with an external domain-specific knowledge, storing professional skills, experiential knowledge, and a comprehensive registry of supported algorithmic workflows. For instance, when the Planner receives a request for an “oppressive atmosphere”, it queries the knowledge base to deconstruct this abstract “vibe" into specific engineering constraints—such as low-key lighting, close-ups, and low-saturation filters—thereby mitigating hallucinations and mediocrity common in general LLMs (Li et al., 2025).
Regarding execution logic, the high-level design adopts a hierarchical orchestration strategy, mapping complex generation tasks through top-down layers of abstraction. The Meta Planner first generates a macro-level SOP blueprint at the creative layer, then propagates this logic to the algorithmic layer to automatically derive and configure the workflow graph structure (Qiu et al., 2025). This hierarchical design ensures the system can both comprehend the “director’s vision” at a macro level and precisely control “technician operations” at a micro level. In essence, the top-level design of Vibe AIGC is not a static toolkit but a dynamic decision flow driven by the Meta Planner: it perceives the user’s “Vibe" in real-time, disambiguates intent via expert knowledge, and ultimately grows a precise, executable workflow from the top down (Xiong et al., 2025).
@@ 5.2 Meta Planner
As the cognitive core of the Vibe AIGC architecture, the Meta Planner assumes the critical responsibility of translating natural language intent into an executable system architecture. Departing from the paradigm of traditional Large Language Models (LLMs) that function merely as text generators or simple routers, the Meta Planner is engineered as a “System Architect" endowed with high-level reasoning capabilities (Goebel and Zips, 2025). Positioned at the forefront of human-computer interaction, it directly interfaces with the user’s natural language input. Its primary function is not the immediate generation of content, but rather deep intent parsing and task decomposition. Upon receiving a user’s “Commander Instructions"—which are often ambiguous and unstructured—the Planner identifies explicit functional requirements while simultaneously capturing latent "Vibe" signals, such as style, mood, and rhythm. Through multi-hop reasoning, it converts these signals into internal logical representations, thereby triggering the entire generative engine.
@@ 5.3 Intent Understanding
The reasoning depth of the Meta Planner stems from its tight synergy with the domain-expert knowledge. This process is designed to flesh out sparse, subjective user instructions into actionable and objective creative schemes, thereby addressing intent sparsity (Fagnoni et al., 2025).
The Planner begins by querying the creative expert knowledge modules, which encapsulate a vast array of multi-disciplinary expertise. For instance, when a user provides the instruction for “a Hitchcockian suspenseful opening," the Planner does not merely treat it as a text prompt. Instead, it leverages the knowledge base to deconstruct the abstract concept of “suspense" into a series of precise creative constraints: visually, it mandates "dolly zoom" camera movements and “high-contrast lighting"; auditorily, it requires "dissonant intervals" in the score; and narratively, it dictates an editing rhythm based on “information asymmetry." Through this process, the Planner externalizes implicit knowledge, transforming the user’s subjective aesthetic intuition into objective, concrete creative scripts. Consequently, this mitigates the issues of “averaging" (mediocrity) or "hallucination" often found in traditional AIGC tools due to comprehension gaps (Shi et al., 2025).